Ballery1:

Find the domain and range of the following function. F(x) = 1/(7-x)(x^2 -1)

brooksn:

Mathway and/or photomath will help you with these questions.

jhonyy9:

- like a first think bc. the denominator of a fraction not can being zero so what will be the restriczions first of ALL ? hope this will help you begining solve it

Ballery1:

Thanks for the suggestion mate.

Narad:

\[f(x)=(x ^{2}-1)/(7-x)\] For the domain, the denominator must be different than zero \[7-x \neq0\] For the range, let \[y=(x ^{2}-1)/(7-x)\] \[y(7-x)=x ^{2}-1\] \[7y-yx=x ^{2}-1\] \[x^{2}+yx-(7y+1)=0\] For this quadratic equation in "x"to have solutions, the discriminant is\[\Delta \ge 0\] \[\Delta = b ^{2}-4ac = y^{2}-4*1*(-(7y+1)) \ge 0\] \[y ^{2}+28y+4 \ge 0\] The solutions to this equations are \[y1=(-28-\sqrt{28^{2}-4*1*4})/2\] and \[y2=(-28+\sqrt{28^{2}-4*1*4})/2\] The range is \[y \epsilon (-\infty , y1] \cup [y2, +\infty )\]